OpenStudy (anonymous):

The area of the sector of a circle with a radius of 8 centimeters is 125.6 square centimeters. The estimated value of π is 3.14. The measure of the angle of the sector is

OpenStudy (anonymous):

A.45 B.135 C.225 D.315 E.450

OpenStudy (moazzam07):

U know full angle of a circle is=360degree

OpenStudy (mathstudent55):

\(\Large A_{sector} = \dfrac{\theta}{360^o} \pi r^2\)

OpenStudy (moazzam07):

Let the angle makes by ur given sector is= theta

OpenStudy (moazzam07):

area of s circle is=pi*r^2

OpenStudy (mathstudent55):

You know A = 125.6 cm^2, r = 8 cm, pi = 3.14. Enter all those values in the formula above and solve for the angle \(\theta\) (Greek letter theta).

OpenStudy (moazzam07):

3.14*4^2=?

OpenStudy (moazzam07):

Required formula is 360/theta=50.24/12.56 Theta=?

OpenStudy (anonymous):

thx for helping out guys

OpenStudy (mathstudent55):

\(\Large A_{sector} = \dfrac{\theta}{360^o} \pi r^2\) \(\Large 125.6~cm^2 = \dfrac{\theta}{360^o} (3.14)(8~cm)^2\) Solve for \(\theta\).

OpenStudy (moazzam07):

np

OpenStudy (mathstudent55):

You're welcome.

OpenStudy (moazzam07):

@mathstudent55 come on man medal me back

OpenStudy (anonymous):

wait my answers arent adding up

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

What did you get?

OpenStudy (anonymous):

i got .625

OpenStudy (mathstudent55):

\(\Large \dfrac{125.6~cm^2(360^o)}{3.14(64 ~cm^2)} = \theta\)

OpenStudy (mathstudent55):

You forgot to multiply by \(360^o\)

OpenStudy (anonymous):

ooooooooooh

OpenStudy (anonymous):

i did it wrong

OpenStudy (anonymous):

thx